# P3Nav: A Unified Framework for Embodied Navigation Integrating Perception, Planning, and Prediction

### v1

@@ 7.1.1 Dataset
For the $\mathrm{CHORES}$-$\mathbb{S}$ ObjectNav benchmark [10], we extend each trajectory with EQA pairs. As a result, we collect 99k EQA pairs as the corresponding EQA dataset for joint training. The $\mathrm{CHORESNAV}$-$\mathbb{S}$ ObjectNavRoom benchmark [10] is similar to the ObjectNav benchmark but involves smaller trajectories. This benchmark includes 15 object categories and annotates 21k trajectories within 2k training houses out of 1M expert trajectory frames. Additionally, the ObjectNavRoom benchmark uses more diverse instructions, describing both the object’s category and its room type simultaneously, such as “Find a vase in the living room.” In contrast, the ObjectNav benchmark specifies only the object’s category, such as “Find a vase.” Similarly, we extend each trajectory in the ObjectNavRoom benchmark with EQA pairs, collecting 21k EQA pairs as the corresponding EQA dataset for ablation studies.
The action space of the ObjectNav and ObjectNavRoom benchmarks [10] includes 20 actions: Move Base (±$\pm 20\mathrm{~{}cm}$); Rotate Base (±$\pm 6^{\circ}$, ±$\pm 30^{\circ}$); Move Arm (x, z) (±$\pm 2\mathrm{~{}cm}$, ±$\pm 10\mathrm{~{}cm}$); Rotate Grasper (±$\pm 10^{\circ}$); pickup; dropoff; done with subtask; and terminate.
@@ 7.2 Traing Strategy
Here, we describe the model hyper-parameters and training details of P3Nav.
@@ 8 Qualitative Results
To visualize the effectiveness of our unified framework for embodied navigation, we provide additional qualitative results generated by our method alongside those of SPOC [10]. As shown in Fig.6, when the agent is positioned close to the target, such as within the same room, our P3Nav is capable of generating the shortest path comparable to SPOC [10]. Both methods understand the environment well, enabling efficient navigation under familiar conditions. However, when the agent needs to navigate across greater distances, such as being situated in different rooms from the target, significant differences in their performance begin to emerge. As shown in Fig. 7, SPOC [10] struggles by repeating paths, which reduces efficiency and increases the risk of looping or missing optimal routes, lowering its success rate. In contrast, P3Nav avoids revisiting areas, systematically explores new routes, and adapts to changing environments, making it effective for long-distance navigation and optimizing complex pathways.
Figure 6: Qualitative comparison of trajectories generated by SPOC [10] and P3Nav in the same room.
Figure 7: Qualitative comparison of trajectories generated by SPOC [10] and P3Nav in different rooms.

### v4

@@ 7.1.1 Dataset
We selected the $\mathrm{CHORES}$-$\mathbb{S}$ benchmark for its complex indoor environments (10K rooms) and diverse object categories (15 types), allowing for comprehensive navigation testing. The $\mathrm{CHORES}$-$\mathbb{S}$ ObjectNav benchmark [10] includes 15 object categories and annotates 99k trajectories within 10k training houses, among 5M expert trajectory frames in the AI2-THOR simulated environment [16]. For the $\mathrm{CHORES}$-$\mathbb{S}$ ObjectNav benchmark [10], we extend each trajectory with EQA pairs. As a result, we collect 99k EQA pairs as the corresponding EQA dataset for joint training. The $\mathrm{CHORESNAV}$-$\mathbb{S}$ ObjectNavRoom benchmark [10] is similar to the ObjectNav benchmark but involves smaller trajectories. This benchmark includes 15 object categories and annotates 21k trajectories within 2k training houses out of 1M expert trajectory frames. Additionally, the ObjectNavRoom benchmark uses more diverse instructions, describing both the object’s category and its room type simultaneously, such as “Find a vase in the living room.” In contrast, the ObjectNav benchmark specifies only the object’s category, such as “Find a vase.” Similarly, we extend each trajectory in the ObjectNavRoom benchmark with EQA pairs, collecting 21k EQA pairs as the corresponding EQA dataset for ablation studies.
The action space of the ObjectNav and ObjectNavRoom benchmarks [10] includes 20 actions: Move Base ($\pm 20\mathrm{~cm}$); Rotate Base ($\pm 6^{\circ}$, $\pm 30^{\circ}$); Move Arm (x, z) ($\pm 2\mathrm{~cm}$, $\pm 10\mathrm{~cm}$); Rotate Grasper ($\pm 10^{\circ}$); pickup; dropoff; done with subtask; and terminate.
@@ 7.2 Traing Strategy
Here, we describe the model hyper-parameters and training details of RoboTron-Nav.
@@ 7.2.2 Training Details
7.2.3 Training Efficiency and Convergence Stability
In terms of training efficiency, multitask training requires approximately twice as much time as single-task training. Additionally, both approaches demonstrate stable loss reduction and typically converge by the fifth epoch.
@@ 8 Qualitative Results
To visualize the effectiveness of our unified framework for embodied navigation, we provide additional qualitative results generated by our method alongside those of SPOC [10]. As shown in Fig.6, when the agent is positioned close to the target, such as within the same room, our RoboTron-Nav is capable of generating the shortest path comparable to SPOC [10]. Both methods understand the environment well, enabling efficient navigation under familiar conditions. However, when the agent needs to navigate across greater distances, such as being situated in different rooms from the target, significant differences in their performance begin to emerge. As shown in Fig. 7, SPOC [10] struggles by repeating paths, which reduces efficiency and increases the risk of looping or missing optimal routes, lowering its success rate. In contrast, RoboTron-Nav avoids revisiting areas, systematically explores new routes, and adapts to changing environments, making it effective for long-distance navigation and optimizing complex pathways.
Figure 6: Qualitative comparison of trajectories generated by SPOC [10] and RoboTron-Nav in the same room.
Figure 7: Qualitative comparison of trajectories generated by SPOC [10] and RoboTron-Nav in different rooms.